OpenStudy (beccaboo333):

Write the equation of the line that passes through (-1, 4) with a slope of -1 in standard form. A. -x + 4y = -1 B. x - 4y = -1 C. x + y = 3 D. x - y = -3

OpenStudy (beccaboo333):

I'm not very great with algebra.. Can I have some help? or maybe the answer.. either one

ganeshie8 (ganeshie8):

(-1, 4) x1 y1 slope = m = -1

ganeshie8 (ganeshie8):

do u knw thepoint-slope form of line ? :) @beccaboo333

OpenStudy (beccaboo333):

nope

ganeshie8 (ganeshie8):

np :) here is the point slope form of line :- y - y1 = m(x - x1)

ganeshie8 (ganeshie8):

just plugin the numbers, wat do u get ?

OpenStudy (beccaboo333):

i still have no idea... i do online classes and usually just wing it at problems.. the pages to learn how to do these problems dont really explain anything well.. so everything is just confusing for me

ganeshie8 (ganeshie8):

i get u, lets try to figure it out quick, its easy actually

ganeshie8 (ganeshie8):

point slope form :- y - y1 = m(x - x1)

ganeshie8 (ganeshie8):

(-1, 4) x1 y1 slope = m = -1

ganeshie8 (ganeshie8):

plug above given data in point slope form :- y-4 = -1(x- -1) y - 4 = -1(x+1)

ganeshie8 (ganeshie8):

can u take it from here ? :)

OpenStudy (beccaboo333):

i can try

ganeshie8 (ganeshie8):

good :) give it a try, if u go wrong, il correct anyways

ganeshie8 (ganeshie8):

just u need to put it in standard form

OpenStudy (beccaboo333):

yeah if i actually was taught this while i was still in public school.. which i never was. so still kinda confusing

ganeshie8 (ganeshie8):

ha it can be confusing when we start, but as u do few problems, im sure u wil feel these are simple

OpenStudy (beccaboo333):

hopefully

ganeshie8 (ganeshie8):

y - 4 = -1(x+1) y - 4 = -x - 1 add x both sides y - 4 = -x - 1 +x +x x + y - 4 = -1

ganeshie8 (ganeshie8):

next, add 4 both sides x + y - 4 = -1 +4 +4 x + y = 3

ganeshie8 (ganeshie8):

thats ur answer see if it makes some sense

OpenStudy (beccaboo333):

it's starting to sorta come back to me how to do this. Can i have some help with 6 more problems? and thank you for the help you've already done :)

ganeshie8 (ganeshie8):

good, please close this question... and ask it as a new question. il try to help if i can :)

OpenStudy (beccaboo333):

how do i close the question?

ganeshie8 (ganeshie8):

scroll to top of this question, u will see "Close quesiton" button. Click it